Dr. Lightman, a psychologist at a private mental hospital was asked to determine whether there was any clear difference in attitudes of smoking for different categories of diagnosis. Each patient was asked about their attitudes toward smoking on a 7-point likert scale (1 = never a good reason for smoking, 7 = always a good reason for smoking). Due to only 7 patients in each category, Dr. Lightman set the alpha value at .10

Diagnosis Category
Emotional Cognitive Drug Related
3 1 6
5 4 7
2 5 4
5 2 4
4 4 6
6 2 5
3 3 3

What are your critical values?
Answer #1

Ans:

we will use ANOVA to determine whether there was any significant difference in attitudes of smoking for different categories of diagnosis.

df(between)=3-1=2

df(within)=21-3=18

alpha=0.1

Critical F value=FINV(0.1,2,18)=2.624